How many ounces of water should a chemist mix with 16 ounces of a 25% sulfuric acid solution to reduce it to a 10% solution?
Found 2 solutions by stanbon, Superownage:
Answer by stanbon(48558)

How many ounces of water should a chemist mix with 16 ounces of a 25% sulfuric acid solution to reduce it to a 10% solution?
-----------------------
25% solution DATA:
Amt = 16 oz ; amt of active ingredient = 0.25*16 = 4 oz
-----------------------
Water solution DATA:
Amt = x; amt of active ingredient = 0*x = 0 oz
------------------------
10% Mixture DATA:
Amt. = 16+x oz ; active ingredient = 0.10(16+x)= 1.6+0.1x oz
-------------------------
EQUATION:
active + active = active
4+0 = 1.6+0.1x
2.4 = 0.1x
x = 24 oz (amount of water needed for the mixture)
================
Cheers,
Stan H.

Answer by Superownage(7)

watch this
25% of 16 is 4,10%=0.10=1/10 the inverse or reciprocal; how to make it an 100% is to multiply by 10,4 x 10 = 40-16=24. your solution is 24.